OpenStudy (anonymous):

Renee took a long flight to visit her family that lives out of the country. She slept for the first half of the trip and watched movies for one-fourth of the remaining flight. If the total trip took ten and one-half hours, how long did Renee watch movies? 1 hour 1 5/16 hours 2 hours 1 and 1/4 hours

OpenStudy (anonymous):

10,5hours = 630 minutes 630/2=315 minutes .. 315 minutes / 4=78,75 minutes 7875/100 minutes / 60 minutes = 1 hours 1875/6000 = 1 hour 5/16 minutes

OpenStudy (anonymous):

thanks :D